### v1

@@ 1 Introduction
In this context, a critical limitation of current segmentation models is their generalizability, particularly their susceptibility to out-of-distribution errors [12]. This issue is exacerbated by the scarcity of annotated medical datasets, which are often prohibitively expensive and time-consuming to compile due to the requirement for expert annotation [13][14]. As a result, many existing models are trained on limited datasets, leading to a tendency to overfit, demonstrating high accuracy on similar data but poor performance when applied to new, unseen images [15]. This lack of resilience when confronted with variations not represented in their training data highlights the pressing need for models that can learn effectively from less labeled data [16][17].
@@
An important aspect of improving segmentation in clinical settings is the consideration of multimodal imaging, where different MRI modalities are utilized to capture unique biological information about a patient [38][39]. Multimodal segmentation tools, which integrate data from various imaging techniques—such as T1-weighted MRI and diffusion-weighted MRI—offer a more comprehensive understanding of complex medical conditions [40]. However, a significant limitation arises when models input all modalities into a single encoder. This approach can constrain the model’s ability to learn modality-specific patterns, which are critical for accurately interpreting the distinct biological targets represented by each modality [41][42]. To address this challenge, we propose a novel modified nnU-Net architecture that incorporates separate encoders for each MRI modality [43]. By learning high-level features independently before merging the knowledge acquired from distinct modalities, our model aims to enhance overall segmentation accuracy, ultimately leading to improved clinical outcomes[44][45].
The challenge of limited labeled data in medical imaging is pervasive, often hindering the development and deployment of effective machine learning models [46][47]. As explained earlier, to mitigate this issue, SSL has gained traction as a viable approach, enabling models to learn from vast amounts of unlabeled data while requiring minimal labeled examples [48][49][50]. SSL techniques can help models develop a foundational understanding of the data, which can then be fine-tuned for specific downstream tasks with limited labeled data [51]. This process involves training the model on a related task or domain and subsequently refining its parameters to adapt to the nuances of the target task, thereby improving performance despite the initial lack of annotated data [52]. Transfer learning also plays a crucial role in this context, allowing models pre-trained on large datasets to be adapted for specific medical imaging challenges [53]. By leveraging existing knowledge, transfer learning can significantly reduce the data requirements for effective model training, enabling more robust and generalizable segmentation outcomes [54][55].
In this paper, we present a comprehensive comparative analysis of various architectural models, including state-of-the-art U-Net models [56] such as vanilla nnU-Net [57] and Multi-encoder nnU-Net [58], versus Transformer models [59]. Additionally, we explore different training strategies, specifically those involving self-supervised learning [60], to assess their impact on model performance [61][62]. By investigating these models and strategies, we aim to elucidate the potential of advanced architectures and learning paradigms in overcoming the current limitations in medical image segmentation, ultimately advancing the state of the art and improving clinical outcomes [63][64].
@@ 2 Related Works
For the 2020 and 2021 BraTS challenges [38], the winning teams proposed the nnU-Net [37], a self-configuring U-Net-based architecture, as a baseline. They implemented several BraTS-specific optimizations, demonstrating its adaptability and effectiveness for tumor segmentation tasks.
@@ 3.1 BraTS Dataset
Imaging Data Description The MRI scans used in the BraTS 2021 challenge consist of four types: a) native (T1), b) post-contrast T1-weighted (T1Gd, using gadolinium), c) T2-weighted (T2), and d) T2 Fluid Attenuated Inversion Recovery (T2-FLAIR) volumes. These scans were acquired using various protocols and scanners from multiple institutions. All BraTS MRI scans underwent standardized pre-processing, which involved converting DICOM files to the NIfTI file format [73], co-registering them to a consistent anatomical template, resampling to a uniform isotropic resolution of 1 mm³, and performing skull stripping. The imaging volumes were segmented using the STAPLE [74] fusion of the top-performing BraTS algorithms, including nnU-Net [37], DeepScan [75], and DeepMedic [41, 42]. These fused labels were manually refined by volunteer neuroradiology experts with varying ranks and experience, adhering to a clearly defined annotation protocol. The final annotations were approved by board-certified attending neuroradiologists with over 15 years of experience in glioma work. The annotated tumor sub-regions are based on known features visible to trained radiologists (VASARI features) and include the Gd-enhancing tumor, peritumoral edematous/invaded tissue, and the necrotic tumor core.
@@ 7 Discussion and Conclusion
In the realm of medical image segmentation, the advent of foundation models, particularly with the integration of SSL, signifies a transformative leap in the precision and efficacy of diagnosing and treating conditions such as tumors [1][2]. The proposed Multi-encoder nnU-Net architecture not only showcases the potential of advanced approaches but also highlights the importance of leveraging multiple MRI modalities to achieve superior segmentation results [37].
@@
Additionally, the challenges of limited labeled data are a pervasive issue in medical imaging, often hindering the development and deployment of effective machine learning models. The two-stage pretraining strategy effectively addresses this limitation by allowing the model to learn from vast amounts of unlabeled data during the self-supervised phase [18, 19]. This innovative approach minimizes the reliance on extensive labeled datasets, which are often prohibitively expensive and time-consuming to compile. The ability to perform well with limited labeled data highlights the model’s robustness and its potential for real-world applications, particularly in clinical settings where annotated data can be scarce.

### v2

@@ 1 Introduction
In this context, a critical limitation of current segmentation models is their generalizability, particularly their susceptibility to out-of-distribution errors [12]. This issue is exacerbated by the scarcity of annotated medical datasets, which are often prohibitively expensive and time-consuming to compile due to the requirement for expert annotation [13][14]. As a result, many existing models are trained on limited datasets, leading to a tendency to overfit, demonstrating high accuracy on similar data but poor performance when applied to new, unseen images [15]. This lack of resilience when confronted with variations not represented in their training data highlights the pressing need for models that can learn effectively from less labeled data [16].
@@
An important aspect of improving segmentation in clinical settings is the consideration of multimodal imaging, where different MRI modalities are utilized to capture unique biological information about a patient [36][37][38]. Multimodal segmentation tools, which integrate data from various imaging techniques—such as T1-weighted MRI and diffusion-weighted MRI—offer a more comprehensive understanding of complex medical conditions [39]. However, a significant limitation arises when models input all modalities into a single encoder. This approach can constrain the model’s ability to learn modality-specific patterns, which are critical for accurately interpreting the distinct biological targets represented by each modality [40][41]. To address this challenge, we propose a novel modified nnU-Net architecture that incorporates separate encoders for each MRI modality. By learning high-level features independently before merging the knowledge acquired from distinct modalities, our model aims to enhance overall segmentation accuracy, ultimately leading to improved clinical outcomes[42][43].
The challenge of limited labeled data in medical imaging is pervasive, often hindering the development and deployment of effective machine learning models [44][45]. As explained earlier, to mitigate this issue, SSL has gained traction as a viable approach, enabling models to learn from vast amounts of unlabeled data while requiring minimal labeled examples [46][47][48]. SSL techniques can help models develop a foundational understanding of the data, which can then be fine-tuned for specific downstream tasks with limited labeled data [17]. This process involves training the model on a related task or domain and subsequently refining its parameters to adapt to the nuances of the target task, thereby improving performance despite the initial lack of annotated data. Transfer learning also plays a crucial role in this context, allowing models pre-trained on large datasets to be adapted for specific medical imaging challenges [49]. By leveraging existing knowledge, transfer learning can significantly reduce the data requirements for effective model training, enabling more robust and generalizable segmentation outcomes [50].
In this paper, we present a comprehensive comparative analysis of various architectural models, including state-of-the-art U-Net models [51] such as vanilla nnU-Net [36] and Multi-encoder nnU-Net [52], versus Transformer models [53]. Additionally, we explore different training strategies, specifically those involving self-supervised learning, to assess their impact on model performance [54][55][56]. By investigating these models and strategies, we aim to elucidate the potential of advanced architectures and learning paradigms in overcoming the current limitations in medical image segmentation, ultimately advancing the state of the art and improving clinical outcomes.
@@ 2 Related Works
For the 2020 and 2021 BraTS challenges [37][58], the winning teams proposed the nnU-Net [36], a self-configuring U-Net-based architecture, as a baseline. They implemented several BraTS-specific optimizations, demonstrating its adaptability and effectiveness for tumor segmentation tasks.
@@ 3.1 BraTS Dataset
Imaging Data Description The MRI scans used in the BraTS 2021 challenge consist of four types: a: native (T1), b: post-contrast T1-weighted (T1Gd, using gadolinium), c: T2-weighted (T2), and d: T2 Fluid Attenuated Inversion Recovery (T2-FLAIR) volumes. These scans were acquired using various protocols and scanners from multiple institutions. All BraTS MRI scans underwent standardized pre-processing, which involved converting DICOM files to the NIfTI file format [74], co-registering them to a consistent anatomical template, resampling to a uniform isotropic resolution of 1 mm³, and performing skull stripping. The imaging volumes were segmented using the STAPLE [75] fusion of the top-performing BraTS algorithms, including nnU-Net [36], DeepScan [46], and DeepMedic [40][41]. These fused labels were manually refined by volunteer neuroradiology experts with varying ranks and experience, adhering to a clearly defined annotation protocol. The final annotations were approved by board-certified attending neuroradiologists with over 15 years of experience in glioma work. The annotated tumor sub-regions are based on known features visible to trained radiologists (VASARI features) and include the Gd-enhancing tumor, peritumoral edematous/invaded tissue, and the necrotic tumor core.
@@ 7 Discussion and Conclusion
In the realm of medical image segmentation, the advent of foundation models, particularly with the integration of SSL, signifies a transformative leap in the precision and efficacy of diagnosing and treating conditions such as tumors [1][2]. The proposed Multi-encoder nnU-Net architecture not only showcases the potential of advanced approaches but also highlights the importance of leveraging multiple MRI modalities to achieve superior segmentation results.
@@
Additionally, the challenges of limited labeled data are a pervasive issue in medical imaging, often hindering the development and deployment of effective machine learning models. The two-stage pretraining strategy effectively addresses this limitation by allowing the model to learn from vast amounts of unlabeled data during the self-supervised phase [17][18]. This innovative approach minimizes the reliance on extensive labeled datasets, which are often prohibitively expensive and time-consuming to compile. The ability to perform well with limited labeled data highlights the model’s robustness and its potential for real-world applications, particularly in clinical settings where annotated data can be scarce.